OpenStudy (anonymous):

CAN SOMEONE CHECK MY ANSWER? The diagonals of kite KITE intersect at point P. If m

OpenStudy (anonymous):

I got A as the answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

Its a kite. x)

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

so, what would be the value of those 2 internal angles?

OpenStudy (anonymous):

you know it's going to be obtuse.... look at it again

OpenStudy (anonymous):

100 DEGREES.

OpenStudy (anonymous):

well close to 90 rather

OpenStudy (anonymous):

THEY EQUAL 180

OpenStudy (anonymous):

<IKE = 180 - 46 - 34 = 100

OpenStudy (anonymous):

you mean 92?

OpenStudy (jdoe0001):

those are equal to 180? they don't look like a flat line

OpenStudy (anonymous):

ohhh..sorry i see that now.

OpenStudy (anonymous):

triangles always = 180

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so you have 2 triangles on the left-hand-side of the rhombus 1 triangle has 46\(^o\) and 90\(^o\) 2 triangle has 34\(^o\) and 90\(^o\)

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so, what about the other 2 angles?

OpenStudy (anonymous):

46^2+90^2= ??? Like that?

OpenStudy (jdoe0001):

ahemm, how many degrees do all INTERNAL angles in a triangle add up to?

OpenStudy (anonymous):

so 46+90+34+90= ??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i cant see the problem here..

OpenStudy (anonymous):

46+90=180 ??

OpenStudy (anonymous):

i dont get it.

OpenStudy (jdoe0001):

how many degrees do all INTERNAL angles in a triangle sum up to you'd say?

OpenStudy (anonymous):

180?

OpenStudy (jdoe0001):

heheh why the "?" character :/

OpenStudy (anonymous):

because i dont know..

OpenStudy (anonymous):

is it 180?

OpenStudy (jdoe0001):

well, my understanding will be that you're supposed to know, otherwise the exercise wouldn't apply

OpenStudy (anonymous):

i dont know. i thought that was the point of this site.. so people who DONT KNOW can get help from people who do know. You know?

OpenStudy (jdoe0001):

so, take a peek of the 2 triangles in the rhombus/kite on the left-hand side you have 2 angles for each what would be the missing angle in each?

OpenStudy (anonymous):

the hypotenuse which would be 180

OpenStudy (jdoe0001):

well, to not understand due to some difference in some pattern, is understandable, not sure about other folks who I see just "doing it for others", but I don't see myself as covering text for others, you're supposed to study the material, if you don't understand it, fine, we can clear it out, now, if you simply skipped it, that's another matter, it's supposed to be covered

OpenStudy (anonymous):

im taking an online class there is no textbook its just quizes. if you cant help me ill go elsewhere. thanks

OpenStudy (anonymous):

Now i have about 5 minutes left for my quiz.. :/

OpenStudy (jdoe0001):

for example for someone with a trigonometry exercise, I don't see myself as showing them how to factor, I do see myself showing to someone doing simple math or arithmetic, but not with trig issues or quadratics

OpenStudy (anonymous):

SOMEONE PLEASE HELP ME! I DONT GET IT.

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Qmb9cnx__hA